An important diamensionless parameters in certain type of fluid flow. problem is froude number define as v/√gl where v= velocity g=acceleration l = length


Determine the froude number for v=10ft/s² g =32.2ft/s² =2ft. Recalculate the froude number using si unit for v,g,and l . Explain the significance of the result of these calculation

1
Expert's answer
2022-02-01T21:26:01-0500

The Froude number is a measurement of bulk flow characteristics such as waves, sand bedforms, flow/depth interactions at a cross-section or between boulders. The denominator represents the speed of a small wave on the water surface relative to the speed of the water, called wave celerity.

It is generally expressed as "Fr = \frac{v}{\sqrt {gd}}"

in which d is a depth of flow, g is the gravitational acceleration (equal to the specific weight of the water divided by its density, in fluid mechanics), v is the celerity of a small surface (or gravity) wave, and Fr is the Froude number.

"Fr = \frac{10}{\sqrt {2\times32.2}} =1.25"

Froude number is greater than 1 for super-critical flow.